OpenStudy (anonymous):

cot x sec^4x = cot x + 2 tan x + tan^3x this is the problem I need to simplify. I'm not certain whether I can think of something like 2 tan as tan^2. I'm also having problems finding identities that will help me. Thank you so much!

OpenStudy (anonymous):

What happens when you sub \(\sec^2(x) = \tan^2(x)+1\)?

OpenStudy (anonymous):

Since \(\sec^4(x) = [\sec^2(x)]^2\)

OpenStudy (anonymous):

So i can use that formula, but instead of tan^2(x) + 1. I do [tan^2(x) + 1]^2?

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

Then foil it out and see if you've gotten closer.

OpenStudy (anonymous):

cot x [tan^2(x) + 1]^2 = cot x + 2 tan x + tan^3x which turns into : cot x tan^4(x^2) + 1 = cot x + 2 tan x + tan^3x correct?

OpenStudy (anonymous):

Not quite... \[ [\tan^2(x)+1]^2 = [\tan^4(x) + 2\tan^2(x) + 1] \]

OpenStudy (anonymous):

Since \(\cot(x) = 1/\tan(x)\) it will basically lower the power of the \(\tan(x)\) terms.

OpenStudy (anonymous):

Sorry, I'm not sure where you got the 2 tan ^2 from